OpenStudy (anonymous):

Divide and simplify: ((x^2-16)/(x+7))/((3x-12)/(x^2-49))

OpenStudy (campbell_st):

start by applying the rule for dividing by fractions, flip the denominator and multiply \[\frac{(x^2 - 16)}{(x + 7)}\times \frac{(x^2 - 49)}{(3x + 12)}\] factor the expressions and you'll see common factors... hope it helps

OpenStudy (anonymous):

Oh is it (x-4)(x-7)/3 ???? @campbell_st

OpenStudy (campbell_st):

yep, that looks good... I'm unsure if you need to distribute for your answer...

OpenStudy (anonymous):

Nope that's one of the answer choices. Thank you!